Q 11 :

Let 9=x1<x2<⋯<x7 be in an A.P. with common difference d. If the standard deviation of x1,x2,…,x7 is 4 and the mean is x¯, then x¯+x6 is equal to          [2023]

  • 18(1+13)

     

  • 2(9+87)

     

  • 34

     

  • 25

     

(3)

Mean =x¯=∑i=17xi7=72[2a+6d]7=a+3d=x4

Variance=∑i=17(xi-x¯)27=(4)2 ⇒∑i=17(xi-x4)27=16

⇒ (3d)2+(2d)2+d2+0+d2+(2d)2+(3d)27=16

⇒ 28d27=16⇒d=2⇒ x¯=9+3(2)=15

x6=a+5d=9+5(2)=19   ∴  x¯+x6=15+19=34



Q 12 :

Let the six numbers a1,a2,a3,a4,a5,a6 be in A.P. and a1+a3=10. If the mean of these six numbers is 192 and their variance is σ2, then 8σ2 is equal to     [2023]

  • 105

     

  • 210

     

  • 200

     

  • 220

     

(2)

Given a1,a2,a3,…,a6 are in A.P.

Sum of these terms in A.P.=62[2a+5d]=192×6

⇒2a+5d=19    ...(i)

Since a1+a3=10⇒a+a+2d=10

⇒2a+2d=10    ...(ii)

Solving (i) and (ii), we get a=2, d=3

So, 2,5,8,11,14,17 are the terms.

Variance, σ2 =22+52+82+112+142+1726-(192)2

=6996-3614=466-3614=1054⇒8σ2=8×1054=210



Q 13 :

The mean and variance of the marks obtained by the students in a test are 10 and 4 respectively. Later, the marks of one of the students is increased from 8 to 12. If the new mean of the marks is 10.2, then their new variance is equal to:            [2023]

  • 3.92

     

  • 3.96

     

  • 4.04

     

  • 4.08

     

(2)

Mean ,μ=10

Variance ,σ2=4

Let the number of observations be n

Sum of observations ∑xi=10n

When marks of one student increased from 8 to 12

∑xi=10n-8+12=10n+4

New mean =10n+4n=10.2⇒n=20; 4=∑xi220-102

Incorrect ∑xi2=104×20=2080⇒Correct ∑xi2=2080-82+122=2160

New σ2=216020-10.22=3.96



Q 14 :

Let S be the set of all values of a1 for which the mean deviation about the mean of 100 consecutive positive integers a1,a2,a3,…,a100 is 25. Then S is         [2023]

  • {99}

     

  • ϕ

     

  • N

     

  • {9}

     

(3)

Let a1 be a natural number.

The 100 consecutive positive numbers are  a1, a1+1, a1+2, a1+3, …, a1+99

Mean (x¯)=a1+a1+1+a1+2+⋯+a1+99100=a1+992

Mean deviation about the mean=∑i=1100|xi2-x¯|100

     =992+972+952+⋯+992100=25

Hence, a1 is a natural number.



Q 15 :

Let the mean and variance of 8 numbers x,y,10,12,6,12,4,8 be 9 and 9.25 respectively. If x>y, then 3x−2y is equal to _____.          [2023]



(25)

We have, x, y, 10, 12, 6, 12, 4, 8

Mean, x¯=x+y+10+12+6+12+4+88=9

⇒x+y=20    ⋯(i)

Now, Variance =x2+y2+(10)2+(12)2+(6)2+(12)2+(4)2+(8)28-81

⇒9.25=x2+y2+5048-81⇒x2+y2=218    ⋯(ii)

Now, (x+y)2=x2+y2+2xy

⇒(20)2=218+2xy⇒2xy=182

Now, (x-y)2=218-182=36⇒x-y=±6    ⋯(iii)

From (i) and (iii), we get x=13,  y=7         (∵x>y)

So, 3x-2y=39-14=25



Q 16 :

Let the mean of the data

x 1 3 5 7 9
Frequency (f) 4 24 28 α 8

 

be 5. If m and σ2 are respectively the mean deviation about the mean and the variance of the data, then 3αm+σ2 is equal to ________ .     [2023]



(8)

Given, x¯=5

⇒∑xifi∑fi=4+72+140+7α+7264+α=5

⇒288+7α=320+5α⇒2α=32⇒α=16

Now, M.D.(x¯)=∑fi|xi-x¯|∑fi

=4×4+24×2+28×0+16×2+8×480=12880=85⇒m=85

Variance (σ2)=∑fi|xi-x¯|2∑fi

=4×16+24×4+28×0+16×4+8×1680=35280=225

∴  3αm+σ2=3×1685+225=486=8



Q 17 :

The mean and standard deviation of the marks of 10 students were found to be 50 and 12 respectively. Later, it was observed that two marks 20 and 25 were wrongly read as 45 and 50 respectively. Then the correct variance is __________.            [2023]



(269)

Given, Mean X¯=50, σ=12

X¯=50-5010=45

∑xi210-502=144                [σ2=∑xi2n-(x¯)2]

∑xi210=2500+144

∑xi210=2500+144+400+625-452-50210=2294

Now correct variance: σT2=2294-(45)2=2294-2025=269



Q 18 :

Let X={11,12,13,…,40,41} and Y={61,62,63,…,90,91} be the two sets of observations. If x¯ and y¯ are their respective means and σ2 is the variance of all the observations in X∪Y, then |x¯+y¯-σ2| is equal to _______ .             [2023]



(603)

Mean x¯=∑i=1141i31=11+412=26

y¯=∑j=6191j31=61+912=76

Combined mean μ=31×26+31×7631+31=26+762=51

σ2=162(∑i=131(xi-μ)2+∑i=131(yi-μ)2)=705

∴ |x¯+y¯-σ2|=|26+76-705|=603



Q 19 :

The mean and variance of 7 observations are 8 and 16 respectively. If one observation 14 is omitted and a and b are respectively mean and variance of remaining 6 observations, then a+3b−5 is equal to ______.              [2023]



(37)

Mean=7,  Variance=16

Let the observations be x1,x2,…,x7.

Now, x1+x2+⋯+x77=8,  x1+x2+⋯+x7=56

When one observation 14 is omitted, the mean is

a=x1+x2+⋯+x7-146⇒a=56-146=7

Variance=∑i=17(xi)27-64=16⇒∑i=17(xi)2=560

When one observation is omitted, then variance

b=∑i=16(xi)26-49⇒b =560-(14)26-49=3646-49=353

So, a+3b-5=7+3×353-5=37



Q 20 :

If the mean and variance of the frequency distribution

xi 2 4 6 8 10 12 14 16
fi 4 4 α 15 8 β 4 5

 

are 9 and 15.08 respectively, then the value of α2+β2-αβ is ________ .              [2023]



(25)

xi fi xifi fixi2
2 4 8 16
4  4 16 64
6 α 6α 36α
8 15 120 960
10 8 80 800
12  β 12β 144β
14 4 56 784
16 5 80 1280
Total 40+α+β 360+6α+12β 3904+36α+144β

 

Mean (x¯)=∑fixi∑fi⇒360+6α+12β40+α+β=9

⇒360+6α+12β=360+9α+9β

⇒3α=3β⇒α=β                               ...(i)

Also, variance σ2=∑fixi2∑fi-(∑fixi∑fi)2

⇒3904+36α+144β40+α+β-(x¯)2=15.08

⇒3904+180α40+2α-(9)2=15.08⇒α=5

From (i), β=5

∴  α2+β2-αβ=25+25-25=25



Q 21 :

If the mean of the frequency distribution

Class : 0–10 10–20 20–30 30–40 40–50
Frequency : 2 3 x 5 4

 

is 28, then its variance is _____ .                                [2023]



(151)

xi fi xi·fi
5 2 10
15 3 45
25 x 25x
35 5 175
45 4 180
  n=14+x ∑xifi=410+25x

 

Mean, x¯=410+25x14+x⇒28=410+25x14+x

⇒392+28x=410+25x⇒3x=18

⇒ x=6

∴  n=20

Variance (σ2)=∑fixi2n-(28)2

=2×52+3×152+6×252+5×352+4×45220-(28)2

=(2×25)+(3×225)+(6×625)+(5×1225)+(4×2025)20-784

=935-784=151



Q 22 :

Let the positive numbers a1,a2,a3,a4 and a5 be in a G.P. Let their mean and variance be 3110 and mn respectively, where m and n are co-prime. If the mean of their reciprocals is 3140 and a3+a4+a5=14, then m+n is equal to _______ .            [2023]



(211)

Let ar2,ar,a,ar,ar2 be positive numbers in G.P.

⇒ ar2+ar+a+ar+ar2=5×3110                         ...(i)

and r2a+ra+1a+1ar+1ar2=5×3140                 ...(ii)

Divide (i) by (ii), we get

a(1r2+1r+1+r+r2)1a(r2+r+1+1r+1r2)= 5×31105×3140

⇒a2=4  

⇒ a=2                                (∵ a cannot be negative)

From (i),     a(1r2+1r+1+r+r2)=5×3110

⇒ (r+1r)2+(r+1r)=314+1

⇒ 4t2+4t-35=0                                                  [Let, r+1r=t]

⇒t=52⇒r=2

∴   Numbers are =12,1,2,4,8

∴   σ2=∑x2N-(∑xN)2=14+1+4+16+645-(3110)2

=34120-961100=18625              ∴ m+n=186+25=211



Q 23 :

If the variance of the frequency distribution

x𝑖 2 3 4 5 6 7 8
Frequency fi 3 6 16 α 9 5 6

 

is 3, then α is equal to _______.                                   [2023]



(5)

xi fi di=xi-5 di2 fidi2 fidi
2 3 - 3 9 27 - 9
3 6 - 2 4 24 - 12
4 16 - 1 1 16 - 16
5 α 0 0 0 0
6 9 1 1 9 9
7 5 2 4 20 10
8 6 3 9 54 18
Total 45+α     150 0

 

σ2=∑fidi2∑fi-(∑fidi∑fi)2=15045+α-(045+α)2

⇒ 3=15045+α    (∵ variance =3)

⇒135+3α=150⇒3α=15⇒α=5



Q 24 :

Let the mean and variance of 8 numbers -10, -7, -1, x, y, 9, 2, 16 be 72 and 2934, respectively. Then the mean of 4 numbers x, y, x+y+1,|x-y| is:             [2026]

  • 9

     

  • 12

     

  • 10

     

  • 11

     

(4)

Mean=-18+x+y+2+9+168=72

⇒x+y+98=72⇒x+y+9=28    .....(1)

Variance=∑zi28-μ2=2934

⇒ 102+72+12+x2+y2+22+92+1628-(72)2=2934    ....(2)

Solving (1) & (2)⇒x=12, y=7

Mean of (1+x+y), x, y,|y-x| is

⇒20+12+7+54=444=11

Option (4)



Q 25 :

Let the mean and variance of 7 observations 2, 4, 10, x, 12, 14, y, x>y, be 8 and 16 respectively. Two numbers are chosen from {1, 2, 3, x-4, y, 5} one after another without replacement, then the probability, that the smaller number among the two chosen numbers is less than 4, is:      [2026]

  • 45

     

  • 13

     

  • 35

     

  • 25

     

(1)

Mean (x¯)=8 (Given)

⇒2+4+10+x+12+14+y7=8

⇒x+y=14    ...(1)

Variance (σ2)=16 (Given)

⇒16=22+42+102+x2+122+142+y27-82

⇒x2+y2=100    ...(2)

∵  (x+y)2=x2+y2+2xy

⇒xy=48   (sum is 14, product is 48)

Since problem states x>y

∴  x=8 and y=6

Now set X={1,2,3,4,6,5}

Now we choose two numbers one after another without replacement 

Total outcomes=6×5=30

We want the probability that the smaller number among the two is less than 4

P(smaller<4)=1-P(smaller≥4)

                             =1-630=45



Q 26 :

The mean and variance of a data of 10 observations are 10 and 2, respectively. If an observation α in this data is replaced by β, then the mean and variance become 10.1 and 1.99, respectively. Then α+β equals:             [2026]

  • 5

     

  • 15

     

  • 10

     

  • 20

     

(4)

Let first 10 numbers are x1,x2,…,x9,α

⇒α+∑i=19xi=100⇒∑i=19xi=100-α

Variance=(∑xi2n)-(∑xin)2

⇒∑xi2n=98

⇒x12+x22+⋯+x92+α2=1020 ⇒∑xi2=1020-α2

In second case, let numbers are x1,x2,…,x9,β

100-α+β=101⇒α-β+1=0

∑xi2+β210-(10.1)2=1.99

⇒β2-α2=20

α=192,  β=212

α+β=19+212=20



Q 27 :

Let X={x∈ℕ:1≤x≤19} and for some a,b∈ℝ,  Y={ax+b:x∈X}. If the mean and variance of the elements of Y are 30 and 750, respectively, then the sum of all possible values of b is                 [2026]

  • 100

     

  • 20

     

  • 80

     

  • 60

     

(4)

∑yi=a∑xi+∑b

=a(1+2+⋯+19)+19b

∑yi19=a·19·202·19+b

30=10a+b  ...(1)

Variance of X=∑xi219-(∑xi19)2

=19·20·3919·6-(10)2=30

Variance of Y=a2(variance of X)

750=a2·30

a2=25⇒a=±5

if a=+5⇒b=30-50=-20  from (1)

if a=-5⇒b=30+50=80  from (1)

sum of values of b=80-20=60

option (4)



Q 28 :

If the mean and the variance of the data

Class

4–8

8–12

12–16

16–20

Frequency

3

λ

4

7

are  μ and 19 respectively, then the value of λ+μ is.  [2026]

  • 18

     

  • 20

     

  • 21

     

  • 19

     

(4)

μ=∑fixi∑fi=18+10λ+56+12614+λ

=200+10λλ+14=10+(60λ+14)

λ+14 is a multiple of 60⇒λ=1,6 or 16

σ2=∑fixi2λ+14-μ2

σ2=62(3)+102(λ)+142(4)+182(7)λ+14-μ2

for λ=6, μ=10+3=13

λ+μ=19



Q 29 :

A random variable X takes values 0, 1, 2, 3 with probabilities 2a+130, 8a-130, 4a+130, b respectively,

where a,b∈R. Let μ and σ respectively be the mean and standard deviation of X such that σ2+μ2=2.

Then ab is equal to :         [2026]

  • 3

     

  • 30

     

  • 60

     

  • 12

     

(3)

x 0 1 2 3
p(x) 2a+130 8a-130 4a+130 b

 

σ2=∑xi2p(xi)-μ2

σ2+μ2=∑xi2p(xi)

=0+1(8a-130)+4(4a+130)+9b

⇒24a+270b+330=2

24a+270b=57

8a+90b=19    ...(1)

Also,

∑p(i)=1

2a+130+8a-130+4a+130+b=1

14a+30b=29    ...(2)

Solving (1) and (2),

a=2,    b=130,    ab=60



Q 30 :

The mean and variance of 10 observations are 9 and 34.2, respectively. If 8 of these observations are 2,3,5,10,11,13,15,21, then the mean deviation about the median of all the 10 observations is   [2026]

  • 6

     

  • 7

     

  • 4

     

  • 5

     

(4)

2+3+5+10+11+13+15+21+a+b10=9

80+a+b10=9⇒a+b=10  ...(1)

∑xi210-(∑xi10)2=34.2

22+32+52+102+112+132+152+212+a2+b210-(9)2=34.2

1094+a2+b210-81=34.2

1094+a2+b2-810=342

a2+b2=58  ...(2)

a=7, b=3  or  a=3, b=7

Numbers =2,3,5,7,10,11,13,15,21

Mean =7+102=8.5

M.D.=6.5+5.5+5.5+3.5+1.5+1.5+2.5+4.5+6.5+12.510

=505

= 5



Q 31 :

The mean and standard deviation of the marks of 200 candidates were found to be 40 and 15 respectively. Later, it was discovered that one student's marks 40 was wrongly read as 50.  The correct mean and standard deviation, respectively are

  • 14.98, 39.95

     

  • 39.95, 14.98

     

  • 39.95, 224.5

     

  • 39.00, 14.00

     

(2)

Corrected ∑x=40×200-50+40=7990

∴ Corrected x¯=7990200=39.95

Incorrect ∑x2=n[σ2+x¯ 2]=200[152+402]=365000

Correct ∑x2=365000-2500+1600=364100

∴ Corrected σ=364100200-(39.95)2=1820.5-1596=224.5=14.98



Q 32 :

An online exam is attempted by 40 candidates out of which 15 are boys. The average marks obtained by boys is 10 with variance 2. The variance of marks obtained by 25 girls is also 2 and the average marks of all 40 candidates is 12.5. If μ is average marks of the 25 girls & σ2 is variance of marks of all 40 candidates then 20σ2−8μ = ______



(3)

μ=40×12.5-10×1525=14

σ2=2+15×2540×40(4)2=5.75



Q 33 :

Let x¯, M and σ2 be respectively the mean, mode and variance of n observations x1,x2,…,xn and di=-xi-a,  i=1,2,…,n where a is any number.

Statement 1:  Variance of d1,d2,…,dn is σ2.

Statement 2:  Mean and mode of d1,d2,…,dn are -x¯-a and -M-a respectively.

  • Statement 1 and Statement 2 are both false

     

  • Statement 1 and Statement 2 are both true

     

  • Statement 1 is true and Statement 2 is false

     

  • Statement 1 is false and Statement 2 is true

     

(2)

S1→V(ax+b)=a2V(x)

S2→mean(ax±b)=a mean(X)±b

Mode(x1,x2,…,xn)=K

Mode(ax1±α,ax2±α,…,axn±α)=ak±α

Central tendencies depend on shift of origin & scaling

Variance depends on scaling but not shift of origin



Q 34 :

∑i=110(xi-5)=5 and ∑i=110(xi-5)2=125, variance of x1,x2,…,x10 will be

  • 94

     

  • 254

     

  • 494

     

  • 814

     

(3)

Let  xi-5=diσx2,  σd2=∑di2n-(∑din)2

=12510-(510)2

=252-14

=494



Q 35 :

The mean and standard deviation of 20 observations are found to be 10 and 2 respectively. On rechecking it was found that an observation by mistake was taken 8 instead of 12. The correct standard deviation is

  • 3.86

     

  • 1.8

     

  • 3.96

     

  • 1.94

     

(3)

∑xi20=10,  ∑xi=10×20=200

If 8 replaced by 12 then ∑xi=200-8+12=204

Now, correct mean (x¯)=∑xi20=20420=10.2

Standard deviation=2

So, variance=(S.D)2=22=4

By definition,  ⇒∑xi220-(∑xi20)2=4⇒∑xi220-(10)2=4

                            ⇒∑xi2=2080

⇒Var=∑xi220-(∑xi20)2=216020-(10.2)2



Q 36 :

The mean and variance of a set of 15 numbers are 12 and 14 respectively. The mean and variance of another set of 15 numbers are 14 and σ2 respectively. If the variance of all the 30 numbers in the two sets is 13, then 3σ2 is equal to

  • 33

     

  • 36

     

  • 30

     

  • 27

     

(3)

∑xi=15×12 and ∑xi215-122=14,    ∑yi=15×14 and ∑yi215-142=σ

Now 13=(14+144)×15+(σ2+196)×1530-132⇒3σ2=30



Q 37 :

If the standard deviation of a1,a2,a3,…,a2025 is λ, then the value of standard deviation of 2a1-1,2a2-1,2a3-1,…,2a2025-1 is

  • 3λ

     

  • 2λ

     

  • λ-1

     

  • 2λ+1

     

(2)

As standard deviation is independent of change of origin

∴  It remains same   ∴ S.D=2λ